# On Calibration of LLM-based Guard Models for Reliable Content Moderation

### v1

@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have facilitated the development of powerful conversation systems, leading to the deployment of LLM-based chatbots in various real-world applications (Brown, 2020; Anil et al., 2023; Touvron et al., 2023; Dubey et al., 2024). However, these systems pose substantial risks due to the potential for malicious exploitation of powerful LLMs by generating harmful content (Wang et al., 2023). Consequently, addressing these risks has become an urgent and critical task. One promising strategy is to regulate LLMs during their training phase. Existing researches primarily focus on designing alignment algorithms through preference optimization (Ouyang et al., 2022; Rafailov et al., 2024), implementing adversarial training (Mazeika et al., 2024), or employing machine unlearning to remove harmful knowledge from the models (Chen & Yang, 2023; Liu et al., 2024). These approaches aim to control text generation and prevent undesired outputs. Despite these significant efforts to enhance LLM safety during training, red-teaming still makes efforts to expose vulnerabilities, including jailbreak attacks that successfully bypass the safety constraint and elicit harmful responses from LLMs, highlighting the risks of future, unseen threats (Zou et al., 2023; Liu et al., 2023; Chao et al., 2024). Therefore, in addition to training-time interventions, it is equally vital to implement test-time measures, such as constraint inference (Xu et al., 2024), and establish effective test-time guardrails through content moderation, particularly when deploying LLMs in real-world settings. Content Moderation serves the critical function of monitoring both user inputs and model outputs during conversations. Typically, guard models are designed to assess whether user inputs and LLM outputs comply with safety regulations, and either reject user queries or block model responses when content violating safety protocols is detected. This approach remains effective even when LLMs have been compromised by previously unseen jailbreak attacks. Current state-of-the-art guard models, which are typically built on LLMs, demonstrate strong performance across various benchmarks (Inan et al., 2023; Ghosh et al., 2024; Han et al., 2024; Zeng et al., 2024). However, these guard models primarily emphasize the classification performance but overlook the predictive uncertainty of harmfulness predictions, therefore failing to assess the reliability of these models’ predictions. This oversight is crucial because guard models may occasionally make erroneous decisions, potentially allowing unsafe content to bypass moderation, especially when encountering non-trivial domain shifts, despite their strong in-domain performances. Therefore, quantifying the predictive uncertainty and confidence in model predictions is essential to assessing the trustworthiness of guard models, enabling more reliable decision-making in high-risk scenarios that may arise during conversations after model deployment.
@@ 2 Related Work
Calibration of LLMs. Confidence calibration is a critical aspect in developing reliable and trustworthy language models (Nguyen & O’Connor, 2015; Guo et al., 2017; Minderer et al., 2021). In the context of LLMs, prior research has explored calibration in question-answering tasks (Jiang et al., 2021) and has empirically examined calibration during the pre-training and alignment stages (Chen et al., 2022; Zhu et al., 2023). Studies such as Lin et al. (2022); Mielke et al. (2022); Xiong et al. (2023) have investigated uncertainty estimation through verbalized confidence, and Kadavath et al. (2022) demonstrated improved calibration of larger models when handling multiple choice and true/false questions given appropriate formats. Another line of research addresses the calibration of biases stemming from in-context samples, instruction templates, sample ordering, and label distribution (Zhao et al., 2021; Zhou et al., 2023b; Liu & Wang, 2023; Fei et al., 2023). These bias calibration techniques indirectly influence the prediction confidence by altering the linear decision boundary (Zhou et al., 2023a), yet they are not designed for explicit confidence calibration. In contrast, our work specifically addresses the challenge of confidence calibration in instruction-tuned guard models for content moderation tasks.
@@ 4.2.1 General Evaluation on public benchmarks
We begin by conducting a comprehensive evaluation of both prompt and response classifications for the existing guard models on public benchmarks. The ECE results for both tasks are presented in Table 1. Our experimental findings indicate that existing guard models exhibit significant miscalibration in both prompt and response classifications. Among the models evaluated, WildGuard demonstrates the lowest average ECE for prompt classification, achieving $14.4\%$, while MD-Judge achieves the lowest average ECE for response classification, at $11.4\%$. However, despite the relatively better performances, both Wildguard and MF-Judge exhibit average ECE values exceeding $10\%$, which is typically considered as poor calibration and underscores the need for further improvements. Additionally, each model displays a substantial variance in ECE across different datasets, suggesting unreliable predictions.
@@ 5.2 Calibration Results
We apply the calibration methods discussed in Section 5.1 to both prompt classification and response classification for each guard model. For temperature scaling, we utilize the XSTest set as the validation set to optimize the temperature due to its relatively small size. This optimized temperature value is then applied across all other datasets, as individual validation sets are not available for all examined datasets. For contextual calibration, we estimate the contextual bias using a space token. For batch calibration, we assume access to the full test set and estimate the contextual bias using the entire test set as a batch. The resulting calibration performance is reported in Table 3. Details regarding the implementation can be found in Appendix A.3, along with additional calibration results in adversarial environments in Appendix A.5.

### v2

@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have facilitated the development of powerful conversation systems, leading to the deployment of LLM-based chatbots in various real-world applications (Brown, 2020; Anil et al., 2023; Touvron et al., 2023; Dubey et al., 2024). However, these systems face substantial risks due to the potential for malicious exploitation of powerful LLMs (Wang et al., 2023). Consequently, addressing these risks has become an urgent and critical task. One promising strategy is to regulate LLMs during their training phase. Existing researches primarily focus on designing alignment algorithms through preference optimization (Ouyang et al., 2022; Rafailov et al., 2024), implementing adversarial training (Mazeika et al., 2024), or employing machine unlearning to remove harmful knowledge from the models (Chen & Yang, 2023; Liu et al., 2024b). These approaches aim to control text generation and prevent undesired outputs. Despite these significant efforts to enhance LLM safety during training, red-teaming still makes efforts to expose vulnerabilities, including jailbreak attacks that successfully bypass the safety constraint and elicit harmful responses from LLMs, highlighting the risks of future, unseen threats (Zou et al., 2023; Liu et al., 2023; Chao et al., 2024; Liu et al., 2024a). Therefore, in addition to training-time interventions, it is equally vital to implement test-time measures, such as constraint inference (Xu et al., 2024), and establish effective test-time guardrails through content moderation, particularly when deploying LLMs in real-world settings.
Content Moderation serves the critical function of monitoring both user inputs and model outputs during conversations. Typically, guard models are designed to assess whether user inputs and LLM outputs comply with safety regulations, and either reject user queries or block model responses when content violating safety protocols is detected. This approach remains effective even when LLMs have been compromised by previously unseen jailbreak attacks. Current state-of-the-art guard models, which are typically built on LLMs, demonstrate strong performance across various benchmarks (Inan et al., 2023; Ghosh et al., 2024; Han et al., 2024; Zeng et al., 2024). However, these guard models primarily focus on the classification performance but overlook the predictive uncertainty of harmfulness predictions (Wang et al., 2024; 2016; Wang & Yeung, 2016; 2020), therefore failing to assess the reliability of these models’ predictions. This oversight is crucial because guard models may occasionally make erroneous decisions, potentially allowing unsafe content to bypass moderation, especially when encountering non-trivial domain shifts (Huang et al., 2022; Xu et al., 2022; Wang et al., 2020), despite their strong in-domain performances. Therefore, quantifying the predictive uncertainty and confidence in model predictions is essential to assessing the trustworthiness of guard models, enabling more reliable decision-making in high-risk scenarios that may arise during conversations after model deployment.
@@ 2 Related Work
Calibration of LLMs. Confidence calibration is a critical aspect in developing reliable and trustworthy language models (Nguyen & O’Connor, 2015; Guo et al., 2017; Minderer et al., 2021). In the context of LLMs, prior research has explored calibration in question-answering tasks (Jiang et al., 2021) and has empirically examined calibration during the pre-training and alignment stages (Chen et al., 2022; Zhu et al., 2023). Studies such as Lin et al. (2022); Mielke et al. (2022); Xiong et al. (2023) have investigated uncertainty estimation through verbalized confidence, and Kadavath et al. (2022) demonstrated improved calibration of larger models when handling multiple choice and true/false questions given appropriate formats. Another line of research addresses the calibration of biases stemming from in-context samples, instruction templates, sample ordering, and label distribution (Zhao et al., 2021; Zhou et al., 2023b; Liu & Wang, 2023; Fei et al., 2023; Abbas et al., 2024). These bias calibration techniques indirectly influence the prediction confidence by altering the linear decision boundary (Zhou et al., 2023a), yet they are not designed for explicit confidence calibration. In contrast, our work specifically addresses the challenge of confidence calibration in instruction-tuned guard models for content moderation tasks.
@@ 4.2.1 General Evaluation on public benchmarks
We begin by conducting a comprehensive evaluation of both prompt and response classifications for the existing guard models on public benchmarks. The ECE results for both tasks are presented in Table 1. Our experimental findings indicate that existing guard models exhibit significant miscalibration in both prompt and response classifications. Among the models evaluated, WildGuard demonstrates the lowest average ECE for prompt classification, achieving $14.4\%$, while MD-Judge achieves the lowest average ECE for response classification, at $11.4\%$. However, despite the relatively better performances, both Wildguard and MF-Judge exhibit average ECE values exceeding $10\%$, which is typically considered a poor calibration and underscores the need for further improvements. Additionally, each model displays a substantial variance in ECE across different datasets, suggesting unreliable predictions.
@@ 5.2 Calibration Results
We apply the calibration methods discussed in Section 5.1 to both prompt classification and response classification for each guard model. For temperature scaling, we utilize the XSTest set as the validation set to optimize the temperature due to its relatively small size. This optimized temperature value is then applied across all other datasets, as individual validation sets are not available for all examined datasets. Additional experiments using in-domain validation sets can be found in Appendix B.1. For contextual calibration, we estimate the contextual bias using a space token. For batch calibration, we assume access to the full test set and estimate the contextual bias using the entire test set as a batch. The resulting calibration performance is reported in Table 3. Details regarding the implementation can be found in Appendix A.3, along with additional calibration results in adversarial environments in Appendix A.5.
@@ 7 Conclusion
Limitations
Our work focuses on post-hoc calibration methods for open-source LLM-based guardrail models. The explored methods do not apply to closed-source models where the logit outputs are unavailable. As for each calibration method, there exist trade-offs. Temperature scaling requires the in-domain validation set for temperature optimization, but in-domain data are not always available in the practical setting. Contextual calibration requires access to the instruction prompt for inference, but the bias captured from content-free tokens may not always be accurate enough. Batch calibration requires access to a batch of unlabeled samples in the target domain, but they could be adapted to adversarial distribution shifts and may need additional validation sets to determine the batch size. In general, post-hoc calibration methods only mitigate the miscalibration in certain scenarios and it is challenging for one single method generalizable to all models and datasets. Nevertheless, this inspires future works to design not only better post-hoc calibration methods but also more reliable training methods to address miscalibration.
@@
Acknowledgments
The authors would like to thank anonymous reviewers for their valuable suggestions. This project is funded by a research grant MOE-MOESOL2021-0005 from the Ministry of Education in Singapore.
@@ A.6 Instruction Prompts
Appendix B Additional Experiments
B.1 Effects of in-domain validation sets for temperature scaling
To further improve the calibration effect of temperature scaling, we conduct additional experiments by optimizing temperature on the in-domain set. Specifically, we consider the WildGuardMix dataset and split a validation set of size 100 from its training set as the in-domain validation set. The temperature is then optimized on the new validation set. The results for prompt and response classification are reported in Table 11. It is observed that the temperature optimized on the in-domain validation set is more effective in reducing ECE than the one optimized on XSTest. Despite the benefits, there is not always access to such an in-domain validation dataset for different scenarios in the wild, leading to a trade-off when applying temperature scaling.
Model Prompt Classification Response Classification Origin TS(XSTest) TS(In-Domain) Origin TS(XSTest) TS(In-Domain) Llama-Guard 28.6 26.0 19.9 14.2 14.0 11.6 Llama-Guard2 26.8 26.0 19.9 12.8 13.6 9.8 Llama-Guard3 20.5 20.4 16.4 11.1 13.0 8.6
Table 11: ECE (%) $\downarrow$ performance comparison of different validation sets.
B.2 Ablation on the length of adversarial prompts
To investigate whether the guard model is vulnerable due to the jailbroken nature of the prompts, or due to some spurious correlations such as length, we design additional experiments. Specifically, we consider three different ranges of adversarial prompt length, leading to three subsets. Then we assess the ECE on three subsets separately and report the results in Table 12. It is observed that there are similar ECE performances among different length ranges, which further supports that it is the jailbroken nature of adversarial prompts that makes guard models vulnerable. We conjecture that there exist many unseen adversarial prompts with any token combinations in the wild and it is impossible to involve all adversarial prompts in instruction tuning of guard models. Thus, the evaluation of adversarial prompts introduces varying levels of uncertainty instead of making overconfident predictions that lead to high ECE performances.
Length (L) Range 0 $\leq$ L $<$ 200 200 $\leq$ L $<$ 500 500 $\leq$ L
Llama-Guard 68.4 76.5 66.2
Llama-Guard2 60.9 61.9 50.7
Llama-Guard3 49.6 46.7 33.4
Table 12: ECE (%) $\downarrow$ performance comparison among different length ranges.
B.3 Robustness on black-box response models
To further assess the ECE performance when classifying responses from proprietary models, we try to collect the responses from GPT-3.5, GPT-4, and Claude-2 to the same set of adversarial queries, and then conduct the same evaluation of ECE. The results are reported in Table 13. It is observed that all guard models have low ECE values when classifying responses from GPT-4 and Claude-2. The reason is that the utilized adversarial attacks are not effective enough for these well-aligned black-box models and thus elicit clear refusal responses that are easy to classify for guard models. Nevertheless, some guard models still exhibit high ECE values when classifying responses from GPT-3.5, serving as a nice complement and support to our Finding 3 in Section 4.2.3.
Guard Model Response Model GPT-3.5 GPT-4 Claude-2 Llama-Guard 39.1 0.0 0.8 Llama-Guard2 18.8 0.0 0.0 Llama-Guard3 0.0 0.0 0.0 Aegis-Guard-D 0.0 0.0 10.7 Aegis-Guard-P 26.6 0.0 0.0 HarmB-Llama 31.2 0.0 0.0 HarmB-Mistral 30.5 0.0 0.0 MD-Judge 21.5 0.0 3.2 WildGuard 9.0 0.0 0.0
Table 13: ECE (%) $\downarrow$ performances of response classification on Harmbench-adv set on black-box response models.
B.4 Comparison of False Positive rate and false negative rates
To investigate what types of over-confidence behaviors different guard models are showing, we report the statistics of the false positive rates (FPR) and false negative rates (FNR) in Table 14. It is observed that most models generally showcase higher false negative rates, suggesting their over-confidence behaviors to predict the input as the safe type.
Model Prompt Classification Response Classification Metric OAI ToxiC SimpST Aegis XST HarmB WildGT BeaverT S-RLHF HarmB WildGT Llama-Guard FPR 8.4 1.4 - 3.2 15.2 - 2.4 8.5 16.6 12.9 0.8 FNR 28.9 53.0 13.0 39.9 17.0 49.8 60.9 46.2 49.5 47.0 67.3 Llama-Guard2 FPR 8.2 3.1 - 4.8 7.6 - 3.6 7.9 21.3 19.3 3.1 FNR 27.4 62.7 8.0 42.5 12.5 11.3 43.6 39.1 28.3 21.5 41.5 Llama-Guard3 FPR 9.2 5.0 - 2.4 2.8 - 4.4 4.0 17.7 35.3 3.5 FNR 21.5 50.0 1.0 43.3 18.0 2.1 34.9 46.0 35.8 3.7 35.6
Table 14: FPR (%) $\downarrow$ and FNR (%) $\downarrow$ performances of prompt and response classification on existing public benchmarks.